Annual Cash Dividend Company per Share $14.00 11.00 9.90 4 1.70 Market Value per Share $ 191.78 123.60 90.83 124.45 Compute tWhich companys stock would probably not be classified as an income stock? Company 1 Company 2 Company 3 Company 4
Answer #1

Dividend yield = Annual dividend / Market price of share

Company 1 = 14/191.78 = 7.30%

Company 2 = 11/123.60 = 8.90%

Company 3 = 9.90/90.83 = 10.90%

Company 4 = 1.70 / 124.45 = 1.37%

Company 4 stock